# Towards Optimal Aggregation of Varying Range Dependencies in Haze Removal

### v2

@@ 3.3.1 Estimating Haze Density Map
where $\mathrm{MSE}()$ is the mean absolute error, and $\mathcal{M}_{d}$ is obtained from $I_{\mathrm{haze}}$ and the learnable paired prompts $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ according to Equation 5. $\alpha_{1}$ and $\alpha_{2}$ are the weights of different loss functions. More details is in supplementary material. Training is lightweight since there is no need to fine-tune the CLIP image encoder.
Finally, we use the learned paired prompts $[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ to generate the estimated patch-wise haze density map $\mathcal{M}_{d}$. Some examples are shown in Figure 5, demonstrating that our method is an effective and universal approach applicable to both homogeneous and non-homogeneous haze.
3.3.2 Obtaining Refined Semantic Map
Considering that the pretrained image encoder of CLIP is obtained through a classification pretext task, each location in the feature map before pooling captures regional semantic information [74]. Therefore, we reuse $F_{\mathrm{image}}$ from Equation 4 and regard it as semantic information. This approach not only provides high-level information but also avoids introducing additional computational overhead.
However, relying solely on $F_{\mathrm{image}}$ leads to suboptimal performance, primarily because it is an abstract representation and lacks low-level semantic details. To address this limitation, we incorporate the input features of the TrambaDA blocks, $F_{\mathrm{input}}$, to supplement $F_{\mathrm{image}}$ and introduce a bidirectional cross-attention mechanism to align their semantic information across different scales, thereby yielding a refined semantic map $\mathcal{M}_{s}$:
$\begin{split}\mathcal{M}_{s}=\mathrm{Linear}&(\text{Attention}_{\text{input}\to\text{image}}(Q_{image},K_{input},V_{input}),\\ &(\text{Attention}_{\text{image}\to\text{input}}(Q_{input},K_{image},V_{image})).\end{split}$ (8)
$Q_{i}$, $K_{i}$ and $V_{i}$ represent the query, key, and value derived from $F_{\mathrm{i}}$ after channel reduction or adaptive pooling, where $\mathrm{i}\in\{\mathrm{image},\mathrm{input}\}$. This process is shown in Figure 4.
3.3.3 Generating Dual-path Aggregation Weights
@@
Here, $\mathrm{Concat}(\cdot)$ denotes a concatenation operation, $\mathrm{Interpolate}(\cdot)$ represents an interpolation operation used to resize $\mathcal{M}_{d}$ and $\mathcal{M}_{s}$ to match the dimensions of $F_{\mathrm{out\_{short}}}$ and $F_{\mathrm{out\_{long}}}$, $\mathrm{Linear}(\cdot)$ refers to a projection layer, and $\mathrm{Split}(\cdot)$ indicates a channel-wise splitting operation. The aggregation is subsequently performed using Equation 3.
4 Experiments
Figure 6: Visual comparisons on the NH-Haze dataset [3]. The red boxes highlight the restoration effects for thick haze and texture details. Please zoom in for the best view.
Figure 7: Visual comparisons on RESIDE SOTS Indoor dataset [38]. The red boxes and green boxes show the restoration of high-frequency and low-frequency information, respectively.
4.1 Experiment Setup
@@
Training Details. DehazeMatic is implemented with PyTorch on NVIDIA A100 GPUs. We use Adam [32] optimizer with default parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$) and a cosine annealing strategy [46] with restarts. The initial learning rate is set to $2\times 10^{-4}$, gradually decreasing to $2\times 10^{-6}$. We train the homogeneous haze dataset for 200 epochs and the non-homogeneous haze dataset for 400 epochs. The images are randomly cropped to a size of 256$\times$256 and augmented with flipping. We use L1 and perceptual loss [29] to supervise dehazing process.
@@
Datasets. We evaluate the proposed DehazeMatic on homogeneous (i.e., RESIDE [38]) and non-homogeneous haze datasets (i.e., NH-HAZE [3] and Dense-Haze [2]). The RESIDE [38] includes two training subsets: the Indoor Training Set (ITS) with 13,990 pairs of indoor images, and the Outdoor Training Set (OTS) with 313,950 pairs of outdoor images. Models are evaluated on the corresponding subset of the Synthetic Objective Testing Set (SOTS). The NH-HAZE [3] and Dense-Haze [2] use a professional haze generator to simulate real hazy conditions, making them suitable for evaluating model performance in real-world non-homogeneous haze removal. These datasets contain 55 and 55 paired images, respectively. We use the last 5 images as the testing set and the remaining images as the training set.
4.2 Comparison with SOTA Methods
Table 1 shows the quantitative results on various benchmarks. As the parameters of the CLIP encoder are frozen, we exclude them from the overhead, which is consistent with previous work. Notably, the additional inference time introduced by CLIP is minimal, requiring only 3.8 ms per image on an NVIDIA A100 GPU.
Performance on Non-homogeneous Haze. MITNet [58] and OKNet [10] are competitive in terms of PSNR and SSIM, respectively; however, our DehazeMatic achieves the highest performance in both metrics across all datasets. Moreover, as shown in Figure 6, MITNet [58] and OKNet [10] suffer from blue artifacts and blurred details, respectively. In contrast, DehazeMatic generates haze-free images with optimal fidelity and richer structural details while maintaining the highest quantitative scores.
Figure 8: Illustration of trade-off between performance and overhead. The line chart in (a) represents the average PSNR on the SOTS-Indoor and SOTS-Outdoor datasets, while the line chart in (b) represents the average PSNR on the NH-Haze and Dense-Haze datasets. The greater the extent to which the line chart’s value exceeds that of the bar chart, the better the model’s trade-off.
@@ 4.2 Comparison with SOTA Methods
Table 2: Quantitative comparisons of various methods on real-world haze dataset RTTS [38]. The best results and the second best results are in bold and underline, respectively.
No-reference Metrics FFA-Net MB-TaylorFormer-B MITNet DehazeFormer DEA-Net OKNet DCMPNet DehazaMatic
@@
ENTROPY $\uparrow$ 7.164 6.907 7.080 7.090 7.079 7.112 7.143 7.405
Performance on Homogeneous Haze. DehazeMatic achieves the SOTA in the average performance of indoor and outdoor scenes, and especially significantly outperforms the second-best method by 0.79 dB with fewer MACs in outdoor scenes. Figure 7 provides visual comparisons on the SOTS Indoor dataset [38]. In the green and red boxed regions, DehazeMatic exhibits superior color consistency and effectively restores finer details, respectively.
Trade-off between Performance and Overhead. As shown in Figure 8, our DehazeMatic achieves the best trade-off for both types of haze, demonstrating its high efficiency.
Generalization on Real-world Haze Dataset. To evaluate generalization, we compare DehazeMatic with several competitive methods on the real-world haze dataset RTTS [38], using models trained on the RESIDE OTS dataset [38] for inference. As shown in Table 2 , our method outperforms all others across no-reference metrics, highlighting its superior generalization ability and practical applicability.
4.3 Ablation Studies
@@
Table 4: Ablation study of CLIP-enhanced Dual-path Aggregator.
@@
Explicit Coexistence of Short- and Long-range Dependencies. We remove either the Mamba path (i.e., long-range) or the Transformer path (i.e., short-range) in the TrambaDA blocks to construct variant models, while simultaneously excluding the CedA module. The corresponding results are shown in the first three rows of Table 3.
Additionally, to verify that its superiority is not dependent on a specific feature extractor, we use convolution [78] (i.e., short-range), linear self-attention [51] (i.e., long-range) and the proposed CedA module to construct a variant model and repeat the above experiments. The corresponding quantitative results are presented in the last three rows of Table 3, demonstrating the versatility of our idea. Although linear self-attention [51] achieves better results than the adopted Mamba block, its time cost is significantly larger, so we do not adopt it. The comprehensive results in Table 3 demonstrate that the complementary short- and long-range dependencies introduced by the dual-path design, along with their reasonable aggregation, can significantly enhance image dehazing.
Figure 9: Visual comparisons of haze density maps estimated using learned prompts and manually predefined prompts.
CLIP-enhanced Dual-path Aggregator (CedA). As shown in Table 4, we conduct extensive ablation studies on the CedA to verify the effectiveness of each component.
@@
(b) We then verify the importance of jointly using haze density and semantic information for aggregation guidance by removing each component individually.
(c) Next, we investigate the estimated haze density map. First, we verify the rationale for using the density map rather than other commonly used guidance signals in dehazing (i.e., the transmission map from DCP [22] and the depth map from Depth Anything [67]) to guide the aggregation. The results indicate that haze density is a more suitable influencing factor.
@@
5 Conclusion
In this paper, we propose DehazeMatic, a framework that employs a dual-path design to explicitly and simultaneously capture complementary local fine-grained details and global contextual cues. Based on key insights from our experiments—namely that different tokens in an image require varying ranges of dependencies and that haze density and semantic information are the primary factors influencing the required range—we introduce the CLIP-enhanced Dual-path Aggregator. It leverages the rich semantic priors and powerful generalization ability encapsulated in CLIP to ensure that each dependency contributes optimally to performance while further exploring CLIP’s potential in image dehazing. Extensive experiments demonstrate that DehazeMatic outperforms SOTA methods on various benchmarks and holds promise for inspiring future research.

### v3

@@ Introduction
where $\mathrm{MSE}()$ is the mean absolute error, and $\mathcal{M}_{d}$ is obtained from $I_{\mathrm{haze}}$ and the learnable paired prompts $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ according to Equation 3. $\alpha_{1}$ and $\alpha_{2}$ are the weights of different loss functions. Training is lightweight since there is no need to fine-tune the CLIP image encoder.
Finally, we use the learned paired prompts $T_{\mathrm{haze}},T_{\mathrm{clear}}]$ to generate the estimated patch-wise haze density map $\mathcal{M}_{d}$. Some examples are shown in Figure 4, demonstrating that our method is an effective and universal approach applicable to both homogeneous and non-homogeneous haze.
Obtaining Refined Semantic Map
Considering that the pretrained image encoder of CLIP is obtained through a classification pretext task, each location in the feature map before pooling captures regional semantic information (Zhang et al. 2022). Therefore, we reuse $F_{\mathrm{image}}$ from Equation 2 and regard it as semantic information. This approach not only provides high-level information but also avoids introducing additional computational overhead.
However, using only $F_{\mathrm{image}}$ gives suboptimal results, primarily because it is an abstract representation and lacks low-level semantic details. To fix this, we incorporate the input features of the TrambaDA blocks, $F_{\mathrm{input}}$, to supplement $F_{\mathrm{image}}$ and introduce a bidirectional cross-attention mechanism to align their semantic information across different scales, thereby yielding a refined semantic map $\mathcal{M}_{s}$:
$\begin{split}\mathcal{M}_{s}=\mathrm{Linear}&(\text{Attention}(Q_{image},K_{input},V_{input}),\\ &(\text{Attention}(Q_{input},K_{image},V_{image})).\end{split}$ (6)
$Q_{i}$, $K_{i}$ and $V_{i}$ represent the query, key, and value derived from $F_{\mathrm{i}}$ after channel reduction or adaptive pooling, where $\mathrm{i}\in\{\mathrm{image},\mathrm{input}\}$. $\text{Attention}(\cdot)$ is self-attention operation.
Generating Dual-path Aggregation Weights
@@
Here, $\mathrm{Concat}(\cdot)$ is concatenation operation, $\mathrm{Interpolate}(\cdot)$ represents an interpolation operation used to resize $\mathcal{M}_{d}$ and $\mathcal{M}_{s}$ to match the dimensions of $F_{\mathrm{out\_{short}}}$ and $F_{\mathrm{out\_{long}}}$, $\mathrm{Linear}(\cdot)$ refers to a projection layer, and $\mathrm{Split}(\cdot)$ indicates a channel-wise splitting operation. The aggregation is subsequently performed using Equation 1.
Experiments
Experiment Setup
@@
Training Details. DehazeMatic is implemented with PyTorch on NVIDIA A100 GPUs. We use Adam (Kingma and Ba 2014) optimizer with default parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$) and a cosine annealing strategy (Loshchilov and Hutter 2016) with restarts. The initial learning rate is set to $2\times 10^{-4}$, gradually decreasing to $2\times 10^{-6}$. We train the homogeneous haze dataset for 200 epochs and the non-homogeneous haze dataset for 400 epochs. The images are randomly cropped to a size of $\times256$ and augmented with flipping. We use L1 loss and perceptual loss (Johnson, Alahi, and Fei-Fei 2016) to supervise dehazing process.
@@
Datasets. We evaluate DehazeMatic on both synthetic and real-world datasets. The synthetic datasets encompass both homogeneous and non-homogeneous haze scenarios. For homogeneous haze, we utilize RESIDE dataset (Li et al. 2018), which includes two training subsets: Indoor Training Set (ITS), containing 13,990 paired indoor images, and Outdoor Training Set (OTS), containing 313,950 paired outdoor images. Evaluation is conducted on the corresponding subsets of the Synthetic Objective Testing Set (SOTS). For non-homogeneous haze, we use the NH-HAZE (Ancuti, Ancuti, and Timofte 2020) and Dense-Haze (Ancuti et al. 2019) datasets, both generated using a professional haze generator to simulate real-world hazy conditions. Each of these datasets comprises 55 paired images, of which the last 5 are used for testing and the remaining 50 for training. For real-world evaluation, we use the RTTS dataset (Li et al. 2018), which contains 4,322 real-world hazy images.
Comparison with SOTA Methods
@@
Table 1: Quantitative comparisons of various methods on multiple synthetic dehazing benchmarks. The best results and the second best results are in bold and underline, respectively.
Methods Homogeneous Haze Non-homogeneous Haze Overhead
SOTS-Outdoor SOTS-Indoor Average NH-Haze Dense-Haze Average
@@
DCP (He, Sun, and Tang 2010) 19.14 0.861 16.61 0.855 17.88 0.858 10.57 0.520 10.06 0.385 10.32 0.453 - -
AOD-Net (Li et al. 2017) 24.14 0.920 20.51 0.816 22.33 0.868 15.40 0.569 13.14 0.414 14.27 0.492 1.76K 0.12G
GridDehazeNet (Liu et al. 2019) 30.86 0.982 32.16 0.984 31.51 0.983 18.33 0.667 14.96 0.533 16.65 0.600 0.96M 21.55G
FFA-Net (Qin et al. 2020) 33.57 0.984 36.39 0.989 34.98 0.987 19.87 0.692 16.09 0.503 17.98 0.598 4.46M 288.86G
DeHamer (Guo et al. 2022) 35.18 0.986 36.63 0.988 35.91 0.987 20.66 0.684 16.62 0.560 18.64 0.622 132.45M 59.25G
SRDefog (Jin et al. 2022) - - - - - - 20.99 0.610 16.67 0.500 18.83 0.555 12.56M 24.18M
MAXIM-2S (Tu et al. 2022) 34.19 0.985 38.11 0.991 36.15 0.988 - - - - - - 14.10M 216.00G
SGID-PFF (Bai et al. 2022) 30.20 0.975 38.52 0.991 34.36 0.983 - - - - - - 13.87M 156.67G
PMNet (Ye et al. 2022) 34.74 0.985 38.41 0.990 36.58 0.988 20.42 0.730 16.79 0.510 18.61 0.620 18.90M 81.13G
MB-TaylorFormer-B (Qiu et al. 2023) 37.42 0.989 40.71 0.992 39.07 0.991 - - 16.66 0.560 - - 2.68M 38.50G
MITNet (Shen et al. 2023) 35.18 0.988 40.23 0.992 37.71 0.990 21.26 0.712 16.97 0.606 19.12 0.659 2.73M 16.42G
DehazeFormer (Song et al. 2023) 34.29 0.983 38.46 0.994 36.38 0.989 20.31 0.761 16.66 0.595 18.49 0.595 4.63M 48.64G
SCANet (Guo et al. 2023) - - - - - - 19.52 0.649 15.35 0.508 17.44 0.579 2.39M 258.63G
DEANet (Chen, He, and Lu 2024) 36.03 0.989 40.20 0.993 38.12 0.991 20.84 0.801 16.73 0.602 18.79 0.702 3.65M 32.23G
UVM-Net (Zheng and Wu 2024) 34.92 0.984 40.17 0.996 37.55 0.990 - - - - - - 19.25M 173.55G
OKNet (Cui, Ren, and Knoll 2024) 35.45 0.992 37.59 0.994 36.52 0.993 20.29 0.800 16.85 0.620 18.57 0.710 4.42M 39.54G
DCMPNet (Zhang, Zhou, and Li 2024) 36.56 0.993 42.18 0.996 39.37 0.995 - - - - - - 18.59M 80.42G
DehazeMatic 38.21 0.995 41.50 0.996 39.86 0.996 21.47 0.806 17.28 0.629 19.38 0.718 4.58M 35.50G
Table 2: Quantitative comparison on the real-world dehazing dataset RTTS. The best results are in bold.
@@
PSD (Chen et al. 2021) 0.920 27.713 4.598
D4 (Yang et al. 2024b) 1.358 33.210 4.484
DGUN (Mou, Wang, and Zhang 2022) 1.111 27.968 4.653
RIDCP (Wu et al. 2023) 0.944 17.293 4.965
CORUN (Fang et al. 2024a) 0.824 11.956 5.342
SGDN (Fang et al. 2025) 0.873 11.549 5.128
DehazeMatic 0.796 11.435 5.510
Figure 5: Visual comparisons on homogeneous hazy images. The bottom displays zoomed-in error maps of selected regions, where darker blue indicates larger restoration errors. Please zoom in to view.
Figure 6: Visual comparisons on non-homogeneous hazy images. The bottom displays zoomed-in error maps of selected regions, where darker blue indicates larger restoration errors. Please zoom in to view.
Figure 7: Illustration of the trade-off between performance and runtime. The blue bar denotes runtime, while the green bar indicates performance. A larger gap, where the green bar exceeds the blue bar, signifies a better trade-off.
Performance on Synthetic Homogeneous Haze
As shown in Table 1, both DCMPNet (Zhang, Zhou, and Li 2024) and MB-TaylorFormer (Qiu et al. 2023) are competitive approaches; however, our DehazeMatic achieves the best overall performance. DCMPNet utilizes depth information as guidance, while MB-TaylorFormer employs linear self-attention for feature extraction. These results demonstrate the superiority of jointly leveraging haze density maps and semantic maps for guidance, as well as simultaneously capturing both short- and long-range dependencies. A visual comparison is shown in Figure 5. Notably, only our method yields minimal restoration errors at object boundaries.
Performance on Synthetic Non-homogeneous Haze
As shown in Table 1, while MITNet (Shen et al. 2023) and OKNet (Cui, Ren, and Knoll 2024) achieve competitive performance in PSNR and SSIM, respectively, our DehazeMatic consistently outperforms them by achieving SOTA results across all metrics. The error maps in Figure 6 demonstrate that our method preserves the most intricate details in the blue-boxed tree region and yields the smallest errors along the object boundaries in the red-boxed target area.
Performance on Real-world Haze
To evaluate the practicality and generalization ability of our model in real-world scenarios, we conduct experiments on the real-world haze dataset RTTS (Li et al. 2018), comparing it with SOTA methods specifically designed for real-world dehazing. As shown in Table 2, our method outperforms all others in terms of no-reference metrics, demonstrating its superiority.
Trade-off between Performance and Runtime
To evaluate the efficiency of our model, we measure its runtime on an NVIDIA A100 GPU and compare the performance and runtime trade-off with the three most competitive methods, as shown in Figure 7. Although our method incorporates CLIP (Radford et al. 2021), the additional inference time it introduces is negligible—only 3.8 ms. As a result, the overall runtime of our model is merely 30.03 ms, achieving the best trade-off. Moreover, our method reaches 33 frames per second (FPS), satisfying real-time processing requirements.
Ablation Studies
@@
Explicit Coexistence of Short- and Long-range Dependencies
To validate the effectiveness of the dual-stream design, we construct variant models by removing either the path responsible for capturing long-range dependencies or the one for short-range dependencies. The corresponding results are presented in the first three rows of Table 3.
Furthermore, to validate the generality of our dual-stream design and show that it does not depend on specific feature extractors in each path, we construct a variant model that employs convolution (Zheng et al. 2023) (i.e., short-range) and linear self-attention (Qiu et al. 2023) (i.e., long-range), and repeat the above experiments. The corresponding quantitative results are reported in the last three rows of Table 3. These results confirm that the integration of complementary short- and long-range dependencies through the dual-stream design substantially improves image dehazing performance.
Table 4: Ablation study of the CedA.
@@
CLIP-enhanced Dual-path Aggregator (CedA)
As shown in Table 4, we conduct extensive ablation studies on the CedA to verify the effectiveness of each component.
@@
(b) We then verify the importance of jointly using haze density and semantic information map for aggregation guidance by removing each component individually.
Figure 8: Visual comparisons of haze density maps estimated by learned prompts and manually predefined prompts.
(c) Next, we investigate the estimated haze density map. First, we verify the rationale for using the density map rather than other commonly used guidance signals in dehazing (i.e., the transmission map from DCP (He, Sun, and Tang 2010) and the depth map from Depth Anything (Yang et al. 2024a)) to guide the aggregation. The results indicate that haze density is a more suitable influencing factor.
@@
Conclusion
We propose DehazeMatic, demonstrating that explicitly capturing complementary short- and long-range dependencies, along with effective aggregation, can achieve SOTA performance and holds potential to inspire future research. To ensure that dependencies across different ranges contribute optimally, we first identify two key factors influencing their relative importance: haze density and semantic information. Guided by this insight, we introduce the CLIP-enhanced Dual-path Aggregator, which generates both maps to instruct the aggregation process. To the best of our knowledge, we are the first to accurately estimate fine-grained haze density maps, further unlocking CLIP’s potential in dehazing.